OpenStudy (anonymous):

ok so figure this out for me: can u figure out the perpendicular slope of J(-10,10), K(7,-9) Please show me the work thanks

OpenStudy (karatechopper):

i can help but i wont solve it for u

OpenStudy (anonymous):

k

OpenStudy (anonymous):

do that then

OpenStudy (karatechopper):

ok

OpenStudy (karatechopper):

ok so what is the slope of those two points?

OpenStudy (anonymous):

-5/12 and 12/-4

OpenStudy (anonymous):

is that rite?: can u help me bra

OpenStudy (karatechopper):

well there is no and..

OpenStudy (karatechopper):

slope=y1-y2/x1-x2

OpenStudy (anonymous):

ok can u just show me how to figure it out for this specific question

OpenStudy (karatechopper):

ok

OpenStudy (karatechopper):

10+9/-10-7 19/-17 thts ur slope^

OpenStudy (karatechopper):

perpendicualr is opposite of reciprocal.. therefore ur slope is now 17/19 now u have equation m=17/19+b

OpenStudy (karatechopper):

now u pick a point